OpenStudy (anonymous):

what is the derivative of the following function y=(x^2+4x+3/√x)

OpenStudy (anonymous):

divide each term by \[\sqrt{x}\] and then use power rule

OpenStudy (anonymous):

how would i devide by each term

OpenStudy (anonymous):

or you can just use the quotient rule

OpenStudy (anonymous):

(2x+4)(x^1/2)-1/2(x^(1/2))/x

OpenStudy (anonymous):

and im sure you can clean it up

OpenStudy (anonymous):

quotient rule is a pain here. think of \[\sqrt{x}=x^{\frac{1}{2}}\] and subtract \[\frac{1}{2}\] from each exponent in the numerator

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

i am assuming it is \[y=\frac{x^2+4x+3}{\sqrt{x}}\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

you would just need to be really strong in your algebra

OpenStudy (anonymous):

then you have \[y=x^{\frac{3}{2}}+4x^{\frac{1}{2}}+3x^{-\frac{1}{2}}\]

OpenStudy (anonymous):

then use power rule and be done

OpenStudy (anonymous):

ok thanks